### v1

# Optimal Allocations under Strongly Pigou-Dalton Criteria: Hidden Layer Structure & Efficient Combinatorial Approach

Abstract.
We investigate optimal social welfare allocations of $m$ items to $n$ agents with binary additive or submodular valuations. For binary additive valuations, we prove that the set of optimal allocations coincides with the set of so-called stable allocations, as long as the employed criterion for evaluating social welfare is strongly Pigou-Dalton (SPD) and symmetric. Many common criteria are SPD and symmetric, such as Nash social welfare, leximax, leximin, Gini index, entropy, and envy sum. We also design efficient algorithms for finding a stable allocation, including an $O(m^{2}n)$ time algorithm for the case of indivisible items, and an $O(m^{2}n^{5})$ time one for the case of divisible items. The first is faster than the existing algorithms or has a simpler analysis. The latter is the first combinatorial algorithm for that problem. It utilizes a hidden layer partition of items and agents admitted by all stable allocations, and cleverly reduces the case of divisible items to the case of indivisible items.
In addition, we show that the profiles of different optimal allocations have a small Chebyshev distance, which is 0 for the case of divisible items under binary additive valuations, and is at most 1 for the case of indivisible items under binary submodular valuations.
Key words and phrases:
Optimal social welfare, Strongly Pigou-Dalton, Binary additive or submodular valuation, Combinatorial algorithm, Layer partition
1. Introduction
Maximizing social welfare or minimizing inequality in allocating resources to agents is an important topic in social economics and has been studied extensively in recent years Coulter (1989); Darmann and Schauer (2015); Atkinson and Bourguignon (2015); Barman et al. (2018). Each agent has her own subjective valuation function over subset of resources (items). Yet how to evaluate the welfare or inequality has no unified answer. Some may suggest LexiMin as the criterion Kleinberg et al. (1999) – where we maximize the smallest valuations of all agents, then maximize the second smallest valuations of them, and so on; while others may suggest Maximum Nash Welfare (MNW) Barman et al. (2018), where we maximize the product of valuations of all agents. For most classes of valuation functions such as the additive valuation functions (where an agent’s valuation is the sum of her valuations for individual items in her bundle), the optimal allocations vary with the selected criteria.
For the special case of binary additive valuations on items, however, Aziz and Rey Aziz and Rey (2021) showed that the LexiMin allocations are equivalent to the MNW allocations. This result raises a natural question that whether there are more connections among other criteria. Benabbou et al. Benabbou et al. (2021) gave a positive answer – the optimal allocations under any strongly Pigou-Dalton (SPD) principle (a.k.a. transfer principle) Moulin (1991) over all allocations with maximal utilitarian social welfare (USW, defined as the sum of valuations, a promise of efficiency) are consistent with LexiMin allocations, even under binary submodular valuations which subsumes binary additive valuations. Roughly, the Pigou-Dalton principle requires that if some income is transferred from a rich person to a poor person, while does not bring the rich to a poorer situation than the poor, then the measured inequality should not increase (or decrease in the strong version); see its formal definition in Section 2 (it is a principle admitted by most common criteria).
Inspired by the aforementioned consistency on optimums among all the SPD criteria Benabbou et al. (2021), we conduct a further study of the optimums under the SPD criteria mainly for the following three scenarios:
IND.:
Indivisible items and agents with binary additive valuations.
DIV.:
Divisible items and agents with binary additive valuations.
IND-SUB.:
Indivisible items and agents with binary submodular valuations.
Our main results are summarized below. We only concern the allocations with maximal USW to ensure the efficiency like Benabbou et al. (2021). Unless otherwise stated, the criteria for evaluating inequality is SPD and symmetric. Moreover, a profile of an allocation $\chi$ refers to the valuations vector of the $n$ agents under $\chi$.
  1. (a)
    For IND, the profiles of optimal allocations have Chebyshev distance at most $1$. Moreover, there is a layer structure hidden behind the optimal allocations – the agents and items are partitioned into serval layers so that items can only be allocated to the agents within the same layer in all optimal allocations.
  2. (b)
    In DIV, the profiles of optimal allocations are all the same (in other words, the Chebyshev distance is 0). (Halpern et al. Halpern et al. (2020) showed a similar result in which only LexiMin and MNW are considered.)
    A layer structure still exists (and with more layers compared to the scenario of IND). More importantly, by utilizing this layer structure and using a reduction to IND, we design the first combinatorial algorithm for finding the optimal allocation for DIV scenario, which runs in $O(m^{2}n^{5})$ time.
  3. (c)
    In IND-SUB, the profiles of optimal allocations have Chebyshev distance at most $1$. This extends our corresponding result in IND. However, the layer structure mentioned above does not hold under this scenario.
Scenario SPD Criteria Consistency Running Time Cheby. Dist. Layer Structure
IND Yes $O(m^{2}n)$ $\leq 1$ Yes
DIV Yes $O(m^{2}n^{5})$ 0 Yes
IND-SUB YesBenabbou et al. (2021) poly.Babaioff et al. (2021) $\leq 1$ No
Table 1. A summary of the results.
See Table 1. In all cases, we derive an “almost consistency” among different optimums (SPD criteria). It states that the valuation of any agent differs by at most $1$ (or 0 for DIV case), under any two different optimal allocations.
Related work.
Halpern et al. Halpern et al. (2020) showed that under binary additive valuations, given any fractional MNW allocation (i.e. the MNW allocations in the scenario of DIV), one can compute, in polynomial time, a randomized allocation with only deterministic MNW allocations (i.e. the MNW allocations in the scenario of IND) in its support and the randomized allocation implements the given fractional MNW allocation. This is a compelling connection between the deterministic and fractional MNW allocations: given a fractional MNW allocation, one can find a convex combination of deterministic MNW allocation to yield it. We note that the connection found by them is not a computational method for fractional MNW allocation (since they need a fractional MNW allocation as input) and our method finds an optimal allocation of the scenario of DIV with algorithm designed for computing optimal allocations of the scenario of IND.
For computational tractability, the SPD optimal allocations can be computed in polynomial time (by computing a Leximin or MNW allocation) in the scenario of IND Kleinberg et al. (1999); Darmann and Schauer (2015); Barman et al. (2018), DIV Nace and Doan (2002) and IND-SUB Aziz and Rey (2021). We propose a method to find an optimal allocation of the scenario of DIV with algorithm designed for computing optimal allocations of the scenario of IND. This is reminiscent of the relation between integer programming and linear programming. The well-known branch-and-bound method uses linear programming as a subprogram to solve integer programming problem. In this paper, based on nontrivial observation, we split each item into a fixed number of pieces and prove that the optimal allocation over the pieces (viewed as indivisible items) is exactly an optimal allocation of the scenario of DIV.
The setting of binary valuation is considered in the resource allocation problem, optimal jobs scheduling, load balancing problem. Lin and Li Lin and Li (2004) studied the special case in which each job can be processed on a subset of allowed machines and its run-time in each of these machines is 1 and find the minimum makespan in polynomial time. Kleinberg et al. Kleinberg et al. (1999) studied case called uniform load balancing which is to assign jobs to machines so that the set of allocated bandwidth is Leximin. The object equals to find the allocation that the number of jobs assigned to machine is Leximax optimal: lexicographically minimizing the number of jobs assigned to machine when sorted from large to small.
Another classic class of resource allocation problems is that with only one kind of resource. In this setting, we only care about the number of items allocated to each agent, rather than the specific subset of items. Ibaraki and Katoh Ibaraki and Katoh (1988) wrote a review on this class of resource allocation problem, viewing as an optimization problem. Allocate a fixed amount of resources (continuous or discrete) to $n$ agents for optimizing the objective function (e.g. separable, convex, minimax, or general). In particular, an important case is the discrete resource allocation problem with a separable convex objective function and Michaeli and Pollatschek Michaeli and Pollatschek (1977) discussed some properties between the optimal solution of the discrete version and the one of continuous version. These properties can be used to speed up the search for integer solution.
The resource problems under ternary valuations, which is a natural extension of binary valuations, are much harder to handle. Under additive valuations, Golovin Golovin (2005) proved that it is NP-hard to compute a $(2-\epsilon)$ approximate maximin allocation even with agents’ valuations of $\{0,1,2\}$ on single items. Another extension of binary valuations is the case where item $j$ has value $p_{j}$ or 0 for each agent. In this case of valuations, for computing maximin allocation, Bezakova and Dani Bezáková and Dani (2005) proved that there is no approximation alogrithm with performance guarantee better than $2$ unless $P=NP$ and Bansal and Sviridenko Bansal and Sviridenko (2006) presented an $O(\frac{\log{\log{m}}}{\log{\log{\log{m}}}})$ approximation algorithm.
2. Preliminaries
For an integer $k>0$, let $[k]$ denote $\{1,...,k\}$. Throughout this paper, $[m]$ refers to the set of $m$ items and $[n]$ refers to the set of $n$ agents.
Each agent $i\in[n]$ has a valuation function $v_{i}:2^{[m]}\rightarrow{\mathbb{R_{+}}}$ over subsets of $[m]$ (called bundles) where $v_{i}(\emptyset)=0$. Given a valuation function $v_{i}$, we define the marginal gain of an item $o$ over a bundle $S\in[m]$ as $\Delta_{i}(S;o)=v_{i}(S\cup{o})-v_{i}(S)$. We focus on the binary valuations where the marginal gain $\Delta_{i}(S;o)\in\{0,1\}$.
Two kinds of binary valuations are discussed frequently in literature, which we call 0/1-add and 0/1-sub. For the 0/1-add valuations, the value of a set of items for an agent is the sum of the valuation of the individual items; the marginal gain $\Delta_{i}(S;o)$ is based on whether agent $i$ likes item $o$ (and independent of $S$). For the 0/1-sub valuations, the marginal gain $\Delta_{i}(S;o)$ does not increase when $S$ grows; formally, $\Delta_{i}(T;o)\leq\Delta_{i}(S;o)$ for $S\subset T\subset[m]$ and $o\in[m]\setminus T$.
Note that the 0/1-sub valuations subsume the 0/1-add ones.
An allocation $\chi$ refers to a collection of disjoint bundles $\chi_{1}\dots\chi_{n}$ such that $\chi_{1}\cup\dots\cup\chi_{n}\subseteq[m]$. An allocation $\chi$ is clean if all the bundles are clean – $\chi_{i}$ is clean if it has no items with zero marginal gain (i.e. for all $o\in\chi_{i}$, $\Delta_{i}(\chi_{i}\setminus\{o\};o)=1$). For 0/1-sub valuations, $\chi_{i}$ is clean if and only if $v_{i}(\chi_{i})=|\chi_{i}|$ (Proposition 3.3 of Benabbou et al. (2021)).
Given a clean allocation $\chi$, assuming agent $i$ gets $h_{i}=|\chi_{i}|$ items under $\chi$, we call vector $(h_{1},\dots,h_{n})$ the profile of $\chi$, denoted by $\mathbf{p}(\chi)$. Henceforth, $h_{i}$ always refers to $|\chi_{i}|$ unless otherwise stated.
Definition 0 (Criterion).
A criterion of income inequality (criterion for short), a.k.a. income inequality metric Atkinson and Bourguignon (2015); Coulter (1989), is a function from the profiles to $\mathbf{R}$: Each profile is evaluated by a real number (called score); the lower the score, the better the profile under this criterion. Following the convention, a criterion must be symmetric; i.e., it should evaluate each permutation of $\mathbf{p}$ the same as $\mathbf{p}$.
Example 0 (Some commonly used criteria).
Let $h^{\uparrow}_{1},\ldots,h^{\uparrow}_{n}$ be the permutation of $h_{1},\ldots,h_{n}$, sorted in increasing order. Take $\Phi(x)$ to be any strictly convex function of $x$. For example, $\Phi(x)=x^{2}$. For every profile $\mathbf{p}=(h_{1},\ldots,h_{n})$, define  [^1] For $\mathsf{NSW}^{\neg}$, we first need to maximize the number of agents with nonzero valuation and then maximize the product of the nonzero valuations.
$\mathsf{NSW}^{\neg}(\mathbf{p})$ $:=-\prod_{h_{i}>0}{h_{i}};$ $\mathsf{Potential}_{\Phi}(\mathbf{p})$ $:=\sum_{i=1}^{n}\Phi(h_{i})$
$\mathsf{GiniIndex}(\mathbf{p})$ $:=\sum_{i}i\cdot h^{\uparrow}_{i};$ $\mathsf{EnvySum}(\mathbf{p})$ $:={\sum}_{h_{i}<h_{j}}(h_{j}-h_{i});$
$\mathsf{Congestion}(\mathbf{p})$ $:=\sum_{i}\binom{h_{i}}{2}$; $\mathsf{Entropy}^{\neg}(\mathbf{p})$ $:=\sum_{i}\frac{h_{i}}{m}\log(\frac{h_{i}}{m}).$
$\mathsf{LexiMax}(\mathbf{p})$ $:=\sum_{i}m^{h_{i}};$ $\mathsf{LexiMin}(\mathbf{p})$ $:=\sum_{i}m^{m-h_{i}};$
Remark 0.
Optimizing $\mathsf{LexiMin}$ is equivalent to maximizing the smallest valuation of all agents, then maximizing the second smallest valuation of them, and so on. Optimizing $\mathsf{LexiMax}$ is equivalent to minimizing the largest valuation of all agents, then minimizing the second largest valuation of them, and so on.
We abbreviate $f(\mathbf{p}(\chi))$ as $f(\chi)$ for any criterion $f$.
Definition 0 (Strongly Pigou-Dalton).
Profile $\mathbf{q}=(q_{1},\ldots,q_{n})$ is regarded more balanced than $\mathbf{p}=(p_{1},\ldots,p_{n})$, if there are $j,k\in[n]$ such that $p_{j}<p_{k}$ and both $q_{j},q_{k}$ lie in $(p_{j},p_{k})$ and $q_{i}=p_{i}$ for $i\in[n]\setminus\{j,k\}$ (namely, the incomes of two agents are more balanced in $\mathbf{q}$ whereas all other incomes remain unchanged). A criterion $f$ is strongly Pigou-Dalton (SPD) Moulin (1991) if $f(\mathbf{q})<f(\mathbf{p})$ whenever $\mathbf{q}$ is more balanced than $\mathbf{p}$. (SPD principle is also known as transfer principle.)
All criteria shown in Example 2 are SPD (proved in Appendix A).
We only consider the allocation with maximal utilitarian social welfare (max-USW, maximizing the sum of the valuations of all agents), otherwise one may minimize $\mathsf{LexiMax}(\chi)$ ($\mathsf{GiniIndex}(\chi)$ etc.) by not allocating any items which is uninteresting. Henceforth, unless otherwise stated, allocations are assumed to be max-USW and clean (we can drop items with zero marginal gain without changing the valuations of agents until the allocation is clean).
3. Indivisible items and agents with 0/1-add valuations
Definition 0 (Stable allocations in IND scenario).
Take an allocation $\chi$ of indivisible items. For each item $o$ allocated to agent $i$ that can be reallocated to another agent $i^{\prime}$ (i.e., $v_{i^{\prime}}(\{o\})=1$), build an edge $(i,i^{\prime})$. Moreover, if there is a simple path $(i_{1},\ldots,i_{k})$ ($k\geq 2$) along such edges, we state that $\chi$ admits a transfer from $i_{1}=u$ to $i_{k}=v$, denoted by $u\rightarrow v$, which consists of k1k-1 reallocations along the path, after which agent $u$ loses and $v$ gains one.
A narrowing transfer refers to a transfer $u\rightarrow v$ with $h_{u}\geq h_{v}+2$. A widening transfer refers to a transfer $u\rightarrow v$ with $h_{u}\leq h_{v}$. Other transfers (i.e. $u\rightarrow v$ with $h_{u}=h_{v}+1$) are called swapping transfers.
Allocation $\chi$ is called nonstable if it admits a narrowing transfer, and is called stable otherwise.
Denote by $\mathcal{S}$ the set of stable allocations.
Lemma 6.
Stable allocations are optimal under $\mathsf{LexiMin}$.
(As a corollary, their profiles are equivalent under permutation.)
See the simple proof of Lemma 6 in Appendix B.  [^2] In fact, Barman et al. Barman et al. (2018) proved that if an allocation is not optimal under $\mathsf{NSW}^{\neg}$, then it admits a narrowing transfer (hence is nonstable). Their result can be easily generalized to any SPD criterion, including $\mathsf{LexiMin}$ (as stated in Lemma 6).
Theorem 7.
1. For any SPD criterion, the optimums are exactly $\mathcal{S}$.
2. We can find a stable allocation in $O(m^{2}n)$ time.
Proof.
1. Fix any SPD criteria $f$, we need to show that
  1. (1)
    A nonstable allocation $\chi$ is non-optimal under $f$.
  2. (2)
    A stable allocation $\chi$ is optimal under $f$.
Together, the optimal allocations under $f$ are the stable ones.
Proof of (1): If $\chi$ is nonstable, it admits a narrowing transfer; denoted by $\chi^{\prime}$ the allocation after this transfer. Clearly, $\mathbf{p}(\chi^{\prime})$ is more balanced than $\mathbf{p}(\chi)$, and therefore $f(\chi^{\prime})<f(\chi)$ due to the assumption that $f$ is strongly Pigou-Dalton.
Proof of (2): Assume $\chi,\chi^{\prime}$ are stable. By Lemma 6, $\mathbf{p}(\chi)$ is equivalent to $\mathbf{p}(\chi^{\prime})$ up to permutation, therefore $f(\chi)=f(\chi^{\prime})$ as $f$ is symmetric (remind that we always assume so). So, stable allocations have the same score under $f$. Further since that nonstable allocations are non-optimal under $f$ (Claim 1), all stable allocations admit the same lowest score under $f$ (a.k.a. optimal).
2. Finding a stable allocation reduces to finding the allocation with minimum $\mathsf{Congestion}$ (by Claim 1 of this theorem), which can be found using network flows. Details are shown in Appendix C.  [^3] The algorithm is almost the same as that of Kleinberg et al. Kleinberg et al. (1999) for finding a $\mathsf{LexiMax}$ optimal allocation but our analysis is simpler.
Remark 0.
It might be asked whether a better allocation under some criterion is also a better allocation under another criterion? This answer is no, although the best remains the best crossing different criteria (Theorem 7). For example, for $m=14$ and $n=3$, we have
$\mathsf{Congestion}((0,5,9))=46<47=\mathsf{Congestion}((2,2,10))$, and
$\mathsf{EnvySum}((0,5,9))=18>16=\mathsf{EnvySum}((2,2,10))$.
3.1. Layer partition of agents and items
Recall that the profiles of stable allocations are equivalent up to permutation (see the corollary part in Lemma 6). Yet the profiles are not unique – e.g., the number of items $h_{i}$ allocated to agent $i$ may differ in different stable allocations. It raises a natural question that to what extent can $\mathbf{p}(\chi)$ differ for different $\chi$ in $\mathcal{S}$?
Our next theorem shows that the difference is negligible:
Theorem 9.
For $\chi,\chi^{\prime}\in\mathcal{S}$, it holds that $|h_{i}-h^{\prime}_{i}|\leq 1$ for each agent $i\in[n]$, where $(h_{1},\ldots,h_{n})=\mathbf{p}(\chi)$ and $(h^{\prime}_{1},\ldots,h^{\prime}_{n})=\mathbf{p}(\chi^{\prime})$. In other words, the Chebyshev distance $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\leq 1$.
Remark 0.
The Chebyshev distance between different optimal allocations remains to be at most 1 even for the scenario of IND-SUB, as stated in Theorem 22 at later part of this paper. The generalized result Theorem 22 is proved in Section 5, and the reader can find a proof of Theorem 9 (which is much easier) in Appendix B.
As the income $h_{i}$ under different stable allocations $\chi$ does not differ too much for every agent $i$, any solution $\chi$ seems to be acceptable for all of them. However, many questions regarding stable allocations remain to be settled. For example:
  • $Q_{1}$.
    Are there common properties of all stable allocations?
  • $Q_{2}$.
    Can we obtain the range of $h_{i}$ under stable allocations?
  • $Q_{3}$.
    How do we count the number of stable allocations?
  • $Q_{4}$.
    Can we find out the profile (of some stable allocation) that optimizes a specific function of $h_{1},\ldots,h_{n}$?
We introduce a “layer partition” of agents and items in the following, which sheds light on the structure of $\mathcal{S}$ and helps answering the above questions.
As a corollary of Theorem 9, in any stable allocation, each agent $i$ falls into two cases: (1) its income $h_{i}$ is a constant $d$ (for all $\chi\in\mathcal{S}$); or (2) its income $h_{i}$ can be $d$ or d1d-1 for some integer $d>0$. Denote
$\mathsf{layer}_{d}$ $=$ $\{i\mid i\text{ falls into the first case}\}$ (1)
$\mathsf{layer}^{-}_{d}$ $=$ $\{i\mid i\text{ falls into the second case}\}.$ (2)
Obviously, $\mathsf{layer}_{0},\mathsf{layer}_{1},\ldots,\mathsf{layer}^{-}_{1},\mathsf{layer}^{-}_{2},\ldots$ form a partition of agents $[n]$, called the partition of agents. Correspondingly, there also exists a partition of items, denoted by $\mathsf{Layer}_{0}$, $\mathsf{Layer}_{1}$, …, $\mathsf{Layer}^{-}_{1}$, $\mathsf{Layer}^{-}_{2}$, …(definition given below), called the partition of items, and the two layer partitions have the following connection:
Under all stable allocations, the items of some layer always belong to those agents in the same layer (to be clear, throughout this paper, $\mathsf{layer}_{d},\mathsf{Layer}_{d}$ are regarded as in the same layer, whereas $\mathsf{layer}^{-}_{d},\mathsf{Layer}^{-}_{d}$ are regarded as in another layer).
The upcoming lemmas demonstrate how do we explicitly define $\mathsf{layer}_{d}$, $\mathsf{layer}^{-}_{d}$, $\mathsf{Layer}_{d}$, $\mathsf{Layer}^{-}_{d}$, (according to which we can compute the entire partition) and also verify the above connection.
Find a fixed stable allocation $\chi$ (using Theorem 7 Claim 2). Recall the transfers on $\chi$ in Definition 5. See Figure 1. Partition $\{i\mid h_{i}=d\}$ into
$p_{d}$ $=$ $\{i\mid h_{i}=d\text{ and }\exists i\rightarrow j\text{ with }h_{j}=d-1\};$ (3)
$q_{d}$ $=$ $\{i\mid h_{i}=d\text{ and }\exists k\rightarrow i\text{ with }h_{k}=d+1\};$ (4)
$r_{d}$ $=$ $\{i\mid h_{i}=d,i\notin p_{d},\text{ and }i\notin q_{d}\}.$ (5)
Figure 1. Computing layer partition of items from $\chi$.
Note that $p_{d}$ is disjoint with $q_{d}$. Otherwise, there is $k\rightarrow j$ with $h_{k}=h_{j}+2$, contradicting the fact that $\chi$ is stable.
For convenience, denote $s_{d}=p_{d}\cup q_{d-1}$. Moreover, denote by $R_{d}$ the set of items allocated to $r_{d}$ under $\chi$, and $S_{d}$ the set of items allocated to $s_{d}$ under $\chi$.
Lemma 11.
Regardless of which stable allocation we choose, those items in $R_{d}$ are always allocated to $r_{d}$, and those items in $S_{d}$ are always allocated to $s_{d}$.
Proof.
For convenience, let sets $r_{0},s_{1},r_{1},s_{2},\ldots$ be indexed increasingly from 0. So as $R_{0},S_{1},R_{1},S_{2},\ldots$. Moreover, assume that the elements in $r_{0},R_{0}$ have index 0; in $s_{1},S_{1}$ have index $1$; in $r_{1},R_{1}$ have index $2$; and etc.
First, observe that (1) a higher-indexed item cannot be allocated to a lower-indexed agent in any (stable) allocation. This follows from two facts (which can be deduced from the definitions of $p_{d},q_{d},r_{d}$ and the assumption that $\chi$ is stable)
Fact 1. there are no edges from $s_{d}$ to $r_{d-1},s_{d-1},r_{d-2},\ldots$, and
Fact 2. there are no edges from $r_{d}$ to $s_{d},r_{d-1},s_{d-1},\ldots$,
It remains to prove that (2) a lower-indexed item cannot be allocated to a higher-indexed agent in any (stable) allocation. We prove this by induction. Here, for the simplicity of presentation, assume $\max\{h_{i}\}=3$ as shown in Figure 1. First, the items indexed lower than $r_{3}$ cannot be allocated to $r_{3}$ in any stable allocation $\chi^{\prime}$. Otherwise, as the items in $R_{3}$ must be allocated to $r_{3}$ (by the analysis above), some agent in $r_{3}$ receives more than 3 items by the pigeonhole principle, which implies that $\mathbf{p}(\chi^{\prime})$ is non-optimal under $\mathsf{LexiMax}$, meaning that $\chi^{\prime}$ is not stable (Theorem 7). By the same argument, (in any stable allocation $\chi^{\prime}$) the items indexed lower than $s_{3}$ cannot be allocated to $s_{3}$ (otherwise $\mathbf{p}(\chi^{\prime})$ contains more agents receiving $3$ items than $\mathbf{p}(\chi)$ which is impossible), the items indexed lower than $r_{2}$ cannot be given to $r_{2}$, so on and so forth. ∎
Lemma 12.
$\mathsf{layer}_{d}=r_{d}$ and $\mathsf{layer}^{-}_{d}=s_{d}$.
Proof.
According to Lemma 11, in every stable allocation, agents $r_{d}$ will receive and only receive $R_{d}$. Note that $|R_{d}|=d|r_{d}|$. Therefore, in every stable allocation, each agent from $r_{d}$ receives exactly $d$ items (otherwise, there must be one agent receiving more than $d$ and one agent receiving less than $d$, which is worse than equally distribution for the criteria $\mathsf{LexiMin}$ and hence nonstable). In other words, the income of an agent from $r_{d}$ is always $d$. Therefore, $r_{d}\subseteq\mathsf{layer}_{d}$.
For each $i\in p_{d}$, we know $h_{i}=d$ and $h_{i}$ can be reduced to d1d-1 (according to (3)), thus $i\in\mathsf{layer}^{-}_{d}$. Therefore $p_{d}\subseteq\mathsf{layer}^{-}_{d}$. For each $i\in q_{d-1}$, we know $h_{i}=d-1$ and $h_{i}$ can be increased to $d$ (according to (4)), thus $i\in\mathsf{layer}^{-}_{d}$. Therefore $q_{d-1}\subseteq\mathsf{layer}^{-}_{d}$. Together, $s_{d}\subseteq\mathsf{layer}^{-}_{d}$.
It follows that $\mathsf{layer}_{d}=r_{d}$ and $\mathsf{layer}^{-}_{d}=s_{d}$. ∎
Combining the last two lemmas, we have $\mathsf{Layer}_{d}=R_{d}$ and $\mathsf{Layer}^{-}_{d}=S_{d}$.
3.2. Some applications of the layer partition – answering $Q_{1},\ldots,Q_{4}$
With the layer partition of agents and items, we can promptly answer $Q_{1}$ to $Q_{4}$.
Answer of $Q_{1}$.:
They all allocate items layer by layer; in which the layers are invariant with respect to the chosen allocation.
Answer of $Q_{2}$.:
For $i\in\mathsf{layer}_{d}$, the range of $h_{i}$ is $[d,d]$. For $i\in\mathsf{layer}^{-}_{d}$, the range of $h_{i}$ is [d1,d][d-1,d]. Computing the layers reduces to computing $r_{d},s_{d}$ for any fixed $\chi$, which is easy. Hence we can compute the ranges of $h_{i}$’s efficiently.
Alternative methods for computing the ranges have to use network flow and are far more complicated.
Answer of $Q_{3}$.:
Counting stable allocations reduces to counting stable allocations within each layer (and then applying the rule of product). However, this counting problem is #P-hard, as the problem of counting stable allocation in $\mathsf{layer}_{1}$ can be reduced to counting perfect matchings which is already #P-hard.
Answer of $Q_{4}$.:
Finding a profile (of a stable allocation) that minimizes a linear function of $h_{1},\ldots,h_{n}$, e.g. $\sum c_{i}h_{i}$, is easy. Modify the network used in the proof of Theorem 7 Claim 2 as follows: Among the edges from $v_{j}$ to $t$, change the cost of the $k$-th one to be $(k-1)\times A+c_{i}$, where $A$ is a large enough constant. Then, the minimum-cost flow still has the minimum $\mathsf{Congestion}$, and it optimizes $\sum c_{i}h_{i}$. For non-linear functions, it is more difficult. Yet the layer partition still helps break down the task.
Remark 0.
Halpern et al. Halpern et al. (2020) implicitly found a structure similar to our layer structures. Specifically, (in the proof of their Theorem 4) given a fractional MNW allocation, they partition the agents into subsets according to the floor of valuations (so in a certain set, the valuation range of each agent is $[x,x+1)$ for some integer $x$). They imply that the agents in each set must be fully allocated to a certain subsets of items for any fractional MNW allocation, and the partition and correspondence between the subsets of agents and the subsets of items also hold for deterministic MNW allocations. We discover our layer structures independently. Our layer structures are more specific, which can be briefly explained through the valuation range of each agent in a certain layer: for the scenario of DIV, compared to their range of $[x,x+1)$ for some integer $x$, the range in our layer structures is some fixed rational number and for the scenario of IND, compared to their range of $[x,x+1]$ for some integer $x$, the range in our layer structures is $[x,x+1]$ for some integer $x$ or some fixed integer (the range of $[x,x]$). Additionally, to compute the layer partition for deterministic MNW allocations, a fractional MNW allocation is necessary through their framework, while we can compute it directly in the scenario of IND (from a deterministic MNW allocation in their term).
4. Divisible items and agents with 0/1-add valuations
This section discusses the scenario of DIV, i.e., the case of divisible items in which we are allowed to allocate a part of an item to an agent (and different parts to different agents perhaps). Note that the 0/1-sub valuations cannot match easily with divisible items. Therefore we restrict ourselves to 0/1-add valuations in this section.
Definition 0 (Stable allocations in scenario of DIV).
Let $\chi$ be an allocation of divisible items. For each part of the item $o$ allocated to agent $i$ that can be reallocated to another agent $i^{\prime}$ (i.e., $v_{i}(\{o\})=1$), build an edge $(i,i^{\prime})$. Moreover, if there is a simple path $(i_{1},\ldots,i_{k})$ ($k\geq 2$) along such edges, we state that $\chi$ admits a transfer from $i_{1}=u$ to $i_{k}=v$, denoted by $u\rightarrow v$. It consists of k1k-1 reallocations with $\Delta>0$ fraction of items along the path, after which $h_{u}$ decreases by $\Delta$ and $h_{v}$ increases by $\Delta$.
A narrowing transfer refers to a transfer from $u$ to $v$ with $h_{u}-\Delta\geq h_{v}+\Delta$. A widening transfer refers to a transfer from $u$ to $v$ with $h_{u}\leq h_{v}$. An Allocation $\chi$ is called nonstable if it admits some narrowing transfer and is stable otherwise.
Denote by $\mathcal{S}^{*}$ the set of stable allocations in DIV scenario.
Lemma 15.
Stable allocations are optimal under $\mathsf{LexiMin}$. (As a corollary, their profiles are equivalent up to permutation.)
Proof.
Assume $\chi$ is non-optimal under $\mathsf{LexiMin}$. We shall prove that $\chi$ is non-stable, i.e., it admits a narrowing transfer.
First, take an allocation $\chi^{*}$ that is optimal under $\mathsf{LexiMin}$.
We build a graph $G$ with $n$ vertices. Be aware that according to $\chi$ and $\chi^{*}$, each item $i$ can be divided into several pieces $i_{1},\ldots,i_{p}$ (with total size $1$), so that each piece is given to a certain agent (denoted by $j$) in $\chi$ and given to a certain agent (denoted by $k$) in $\chi^{\prime}$. If $j\neq k$, build an arc from $j$ to $k$, with weight equal to the size of this piece. Clearly, an arc represents a reallocation (of one piece) on $\chi$, and $\chi$ becomes $\chi^{*}$ after all the arcs (i.e. reallocations) are applied.
We decompose graph $G$ into several cycles $C_{1},\ldots,C_{a}$ and paths $P_{1},\ldots,P_{b}$, where the edges in any cycle or path have the same weight, and where $t_{j}\neq s_{k}$ for $j\neq k$, where $s_{i},t_{i}$ denote the starting and ending vertices of $P_{i}$, respectively. Such a decomposition exists under appropriate division of items.
For $0\leq i\leq b$, let $\chi^{(i)}$ be the allocation copied from $\chi$ but applied all the arcs (reallocations) in $P_{1},\ldots,P_{i}$. $\chi^{(0)}=\chi$.
Be aware that $\chi^{(b)}$ becomes $\chi^{*}$ after applying the arcs in $C_{1},\ldots,C_{a}$. We obtain that $\mathsf{LexiMin}(\chi^{(b)})=\mathsf{LexiMin}(\chi^{*})$. Further since that $\mathsf{LexiMin}(\chi^{*})<\mathsf{LexiMin}(\chi)$, there exists $i\penalty 10000\ (1\leq i\leq b)$ such that $\mathsf{LexiMin}(\chi^{(i)})<\mathsf{LexiMin}(\chi^{(i-1)})$.
It follows that in $\chi^{(i-1)}$, we have $h_{s}>h_{t}$ (where $s,t$ denote $s_{i},t_{i}$ respectively, for short). It further follows that in $\chi^{(0)}=\chi$, we also have $h_{s}>h_{t}$, since $h_{s}$ never increases and $h_{t}$ never decreases in the sequence $\chi^{(0)},\ldots,\chi^{(i-1)}$. Consequently, $\chi$ admits a narrowing transfer (from $s$ to $t$). ∎
Theorem 16.
1. For any SPD criterion, the optimums coincide with $\mathcal{S}^{*}$. 2. We can find a stable allocation in polynomial time.
Just like we prove Theorem 7 Claim 1 using Lemma 6, we can prove Theorem 16 Claim 1 using Lemma 15 (the proof is omitted).
To find a stable allocation in polynomial time, we can use an approach based on linear programming (LP) which is shown in Section 4.2 (thus we prove Theorem 16 Claim 2). Alternatively, we find a purely combinatorial approach (which is more efficient and more interesting) for finding such an allocation. It utilizes the layer partition with several nontrivial ideas. See details in Section 4.1.
Theorem 17.
For $\chi,\chi^{\prime}\in\mathcal{S}^{*}$, it holds that $\mathbf{p}(\chi)=\mathbf{p}(\chi^{\prime})$, namely, the Chebyshev distance $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))=0$.
A proof of Theorem 17 is related to the LP approach mentioned above and is given in Section 4.2. An alternative proof is similar to the proof of Theorem 9 and is omitted.
4.1. Layer partition of agents and items for divisible items & a combinatorial algorithm for finding a stable allocation
In the following, we extend the layer partition given in Section 3.1 to the divisible case and then present the aforementioned combinatorial algorithm.
According to Theorem 17, profile $\mathbf{p}(\chi)$ is unique for $\chi\in\mathcal{S}^{*}$. In other words, the income $h_{i}$ of each agent $i$ is independent of $\chi$, as long as $\chi$ is stable.
For each real number $d$, denote by $\mathsf{layer}^{*}_{d}$ the set of agents that always receive $d$ items no matter in which stable allocation; formally, $\mathsf{layer}^{*}_{d}=\{i\mid h_{i}=d\}$.
Lemma 18.
The set of items allocated to $\mathsf{layer}^{*}_{d}$ is invariant for $\chi\in\mathcal{S}^{*}$. Moreover, this set (denoted by $\mathsf{Layer}^{*}_{d}$ henceforth) consists of complete items only.
Proof.
Fix $\chi$. Let $L_{d}$ be the items allocated to $\mathsf{layer}_{d}$ under $\chi$. An item in $L_{d^{\prime}}$ ($d^{\prime}>d$) cannot be allocated to $\mathsf{layer}_{d}$ otherwise there is a narrowing transfer in $\chi$. An item in $L_{d^{\prime}}$ ($d^{\prime}<d$) cannot be allocated to $\mathsf{layer}_{d}$ otherwise the allocation is not $\mathsf{LexiMax}$ optimal (can be proved by induction as in the proof of Lemma 11). Therefore, those items in $L_{d}$ can only be allocated to $\mathsf{layer}_{d}$ (even for other stable allocations). We thus obtain the first part of this lemma.
2. In any stable allocation, an item cannot be allocated to different layers. Otherwise there is clearly a simple narrowing transfer. ∎
Lemma 18 implies a layer partition of items and agents, where items $\mathsf{Layer}^{*}_{d}$ and agents $\mathsf{layer}^{*}_{d}$ are in the same layer.
Lemma 19.
A stable allocation $\gamma$ for the divisible case can be obtained by optimally reallocating items $\mathsf{Layer}^{-}_{d}$ (regarded as divisible items) to the agents $\mathsf{layer}^{-}_{d}$ and items $\mathsf{Layer}_{d}$ (regarded as indivisible items) to $\mathsf{layer}_{d}$ for each $d$.
Proof.
For each $d$, let $\gamma^{-}_{d}$ be an optimal allocation of items $\mathsf{Layer}^{-}_{d}$ (regarded as divisible items) to the agents $\mathsf{layer}^{-}_{d}$. Moreover, let $\gamma_{d}$ be an optimal allocation of items $\mathsf{Layer}_{d}$ (regarded as indivisible items) to $\mathsf{layer}_{d}$ so that each agent in $\mathsf{layer}_{d}$ receives $d$ items. Combine $\gamma_{d},\gamma^{-}_{d}(d\geq 0)$ in all layers to obtain an overall allocation $\gamma$. We claim that $\gamma\in\mathcal{S}^{*}$. The proof is as follows.
Those items in higher layer cannot be given to agents in lower layer – there are no such edges as shown in the proof of Lemma 11 (this also holds for the divisible case). Hence there is no narrowing transfer between layers in $\gamma$. Also, there is no narrowing transfer within any layer of $\gamma$ by the construction of $\gamma$. Together, $\gamma$ admits no narrowing transfer and hence is stable. So, $\gamma\in\mathcal{S}^{*}$. ∎
Lemma 20.
For any $\chi\in\mathcal{S}^{*}$, it holds that
1. $h_{i}=d$ for each $i\in\mathsf{layer}_{d}$. 2. $h_{i}\in[d-1,d]$ for $i\in\mathsf{layer}^{-}_{d}$.
Proof.
Denote $(g_{1},\ldots,g_{n})=\mathbf{p}(\gamma)$ and $(h_{1},\ldots,h_{n})=\mathbf{p}(\chi)$.
Combining (1)-(3) below, we immediately obtain the lemma.
  • (1)
    For each $i\in[n]$, it holds that $h_{i}=g_{i}$ (apply Theorem 17 with $\gamma,\chi\in\mathcal{S}^{*}$).
  • (2)
    For each $i\in\mathsf{layer}_{d}$, it holds that $g_{i}=d$ (trivial).
  • (3)
    For each $i\in\mathsf{layer}^{-}_{d}$, it holds that $g_{i}\in\left[d-1,d\right]$. (Proof: Since $\gamma^{-}_{d}$ is ($\mathsf{LexiMin}$) optimal, $g_{i}\geq d-1$. Since $\gamma^{-}_{d}$ is also ($\mathsf{LexiMax}$) optimal, $g_{i}\leq d$.)
Lemma 21.
Given $\chi\in\mathcal{S}^{*}$, for two different layers $\mathsf{layer}^{*}_{d}$ and $\mathsf{layer}^{*}_{d^{\prime}}$ of $\chi$, we have
$\left|d-d^{\prime}\right|\geq\frac{1}{n^{2}}.$
Proof.
By Lemma 19, we obtain that
$d=\frac{|\mathsf{Layer}^{*}_{d}|}{|\mathsf{layer}^{*}_{d}|},\quad d^{\prime}=\frac{|\mathsf{Layer}^{*}_{d^{\prime}}|}{|\mathsf{layer}^{*}_{d^{\prime}}|}.$
Recall Lemma 18. $|\mathsf{Layer}^{*}_{d}|$ and $|\mathsf{Layer}^{*}_{d^{\prime}}|$ are integers. Thus,
$\left|d-d^{\prime}\right|=\left|\frac{|\mathsf{Layer}^{*}_{d}||\mathsf{layer}^{*}_{d^{\prime}}|-|\mathsf{Layer}^{*}_{d^{\prime}}||\mathsf{layer}^{*}_{d}|}{|\mathsf{layer}^{*}_{d}||\mathsf{layer}^{*}_{d^{\prime}}|}\right|\geq\frac{1}{n^{2}}.$
The last three lemmas are crucial to the combinatorial algorithm.
For divisible case, we divide each item into $2n^{2}$ identical (unbreakable) pieces with size $\frac{1}{2n^{2}}$, and solve this indivisible case (with $2n^{2}m$ indivisible pieces) using the algorithm in Theorem 7 Claim 2.
We obtain an allocation with several layers, each layer’s index $d$ is a real number – a multiple of $\frac{1}{2n^{2}}$. For those agent $j\in\mathsf{layer}_{d}$, we have $h_{j}=d$, and for those $i\in\mathsf{layer}^{-}_{d}$, we have $h_{i}\in[d-\frac{1}{2n^{2}},d]$.
According to Lemma 20, for the divisible case, we have
$h_{i}\in[d-\frac{1}{2n^{2}},d]\text{ for }i\in\mathsf{layer}^{-}_{d}.$
As a result, $i\in\mathsf{layer}^{*}_{d^{\prime}}$, for some $d^{\prime}\in[d-\frac{1}{2n^{2}},d]$.
By Lemma 21, there can only be one such $d^{\prime}$ in the range of $d-\frac{1}{2n^{2}},d]$ for $i\in\mathsf{layer}^{-}_{d}$. By Lemma 19, we can obtain an optimal allocation for divisible case just by reallocating items within each layer, and we can calculate that $d^{\prime}=\frac{|\mathsf{Layer}^{-}_{d}|}{|\mathsf{layer}^{-}_{d}|}.$
The obtained algorithm deals with $2mn^{2}$ items and $n$ agents. Therefore, it has a time complexity of $O(m^{2}n^{5})$, which is still better than the linear programming procedure (explained below) with time complexity of $O(m^{3.5}n^{5.5})$ with $O(mn)$ variables and $O(n^{2})$ linear programming problems in the worst case.
4.2. Find a stable allocation for divisible items using LP and a proof of Theorem 17
This subsection provides an alternative approach based on linear programming for finding a stable allocation for divisible items. This approach is more straightforward but it is not combinatorial.
Applying Theorem 16 Claim 1, finding a stable allocation reduces to finding an allocation with the minimum $\mathsf{LexiMin}$. We claim that
  • (1)
    the latter further reduces to computing the “fair multi-flow” Nace and Doan (2002) in the network below:
Figure 2. Reduction to fair multi-flow problem.
See Figure 2. There are $n$ source nodes $s^{1},\ldots,s^{n}$ in the first layer $S$. The second layer $R$ consists of only one node $r$ as a relay. The third layer $I$ has $m$ nodes, standing for items. The last layer $T$ has $n$ sink nodes $t^{1},\ldots,t^{n}$. For each agent $d$, there is an arc $(s^{d},r)$ of unlimited capacity. For every item $i$ in $I$, there is an arc $(r,i)$ of capacity 1. If item $i$ can be allocated to agent $d$, an arc $(i,t^{d})$ is added with capacity 1.
In the aforementioned fair multi-flow problem, we shall send a flow $f_{d}$ from $s^{d}$ to $t^{d}$ for each $d$. The sum of the flows on each edge cannot exceed the capacity, and the $n$ amount of flows $|f_{1}|,\ldots,|f_{n}|$ as a vector should have the lowest $\mathsf{LexiMin}$. Clearly, the solution to this problem corresponds to the optimal allocation under $\mathsf{LexiMin}$ (items are fully allocated as the multi-flow is $\mathsf{LexiMin}$), therefore claim (1) holds.
Finally, recall the result of Nace and Doan (2002) which states that the fair multi-flow problem can be reduced to a polynomial number of linear programming problems and thus be solved in polynomial time.
Proof of Theorem 17.
Recall finding a stable allocation by computing a fair multi-flow, the flow vector $(|f_{1}|,\ldots,|f_{n}|$) is unique for multi flow problem, according to Nace and Doan (2002). Therefore the profile of optimal allocation is also unique. ∎
5. Indivisible items and agents with 0/1-sub valuations
We now move on to the scenario of IND-SUB.
The 0/1-sub function is closely related to the matroid theory Oxley (2011). A matroid is a pair $(E,\mathcal{I})$, where $E$ is a finite set (called the ground set) and $\mathcal{I}$ is a family of subsets of $E$ (called the independent sets).
The independent sets satisfy the following three axioms:
  1. (I1)
    $\emptyset\in\mathcal{I}$,
  2. (I2)
    if $Y\in\mathcal{I}$ and $X\subseteq Y$, then $X\in\mathcal{I}$, and
  3. (I3)
    if $X,Y\in\mathcal{I}$ and $|X|<|Y|$, then there exists $y\in Y\setminus X$ such that $X\cup\{y\}\in\mathcal{I}$.
If an agent has a 0/1-sub valuation on items, then the set of clean bundles forms the set of independent sets of a matroid (proved in Benabbou et al. Benabbou et al. (2021)). Benabbou et al. Benabbou et al. (2021) further showed that under 0/1-sub valuations, the set of max-USW allocations (under any SPD criterion) are consistent with $\mathsf{LexiMin}$ allocations. This result generalizes Theorem 7 Claim 1 (for the 0,1-add valuations). Besides, Babaioff et al.Babaioff et al. (2021) showed that under 0/1-sub valuation, the allocation that optimizes $\mathsf{NSW}^{\neg}$ can be found in polynomial time.
To the best of our knowledge, however, it is not known to what extent can $\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime})$ differ for different optimal allocations $\chi,\chi^{\prime}$. We answer this question by the following theorem.
Let $\mathcal{SS}$ denote the set of (max-USW and clean) allocations which are optimal under any SPD criterion.
Theorem 22.
For $\chi,\chi^{\prime}\in\mathcal{SS}$, it holds that $|h_{i}-h^{\prime}_{i}|\leq 1$ for each agent $i\in[n]$, where $(h_{1},\ldots,h_{n})=\mathbf{p}(\chi)$ and $(h^{\prime}_{1},\ldots,h^{\prime}_{n})=\mathbf{p}(\chi^{\prime})$. In other words, the Chebyshev distance $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\leq 1$.
Proof.
Assume (i) $\chi,\chi^{\prime}\in\mathcal{SS}$ and (ii) $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\geq 2$. If there are multiple such pairs of allocations, take the pair $(\chi,\chi^{\prime})$ with minimum symmetric difference $\sum_{i\in[n]}|\chi_{i}\triangle\chi^{\prime}_{i}|$.
Without loss of generality, assume that
$h_{1}\leq h_{2}\leq\dots\leq h_{n}\text{ and }h_{q}\geq h^{\prime}_{q}+2.$
Assume that $h^{\prime}_{j_{1}}\leq h^{\prime}_{j_{2}}\leq\dots\leq h^{\prime}_{j_{n}}$.
Notice that $\mathbf{p}(\chi)\neq\mathbf{p}(\chi^{\prime})$. Take the minimum index $i$ satisfying $h_{i}\neq h^{\prime}_{i}$. Clearly, for all $k\in[i-1]$ we have
$h_{k}=h^{\prime}_{k}$ (6)
In fact, it must hold that
$h_{i}<h^{\prime}_{i}$ (7)
Recall that the smaller $\mathsf{LexiMin}(\mathbf{p})$, the better the vector $\mathbf{p}$ under $\mathsf{LexiMin}$. Indeed, if $h_{i}>h^{\prime}_{i}$, then together with (6),
$\mathsf{LexiMin}((h_{1},\dots,h_{i}))>\mathsf{LexiMin}((h^{\prime}_{1},\dots,h^{\prime}_{i})).$
Moreover,
$\mathsf{LexiMin}((h^{\prime}_{1},\dots,h^{\prime}_{i}))\geq\mathsf{LexiMin}((h^{\prime}_{j_{1}},\dots,h^{\prime}_{j_{i}}))$
since $h^{\prime}_{j_{1}},\dots,h^{\prime}_{j_{i}}$ are the smallest $i$ elements in $\mathbf{p}(\chi^{\prime})$. Together, we have $\mathsf{LexiMin}((h_{1},\dots,h_{i}))>\mathsf{LexiMin}((h^{\prime}_{j_{1}},\dots,h^{\prime}_{j_{i}}))$. By comparing the smallest $i$ elements of $\mathbf{p}(\chi)$ and $\mathbf{p}(\chi^{\prime})$, we see $\mathsf{LexiMin}(\mathbf{p}(\chi))>\mathsf{LexiMin}(\mathbf{p}(\chi^{\prime}))$, contradicting the assumption that $\chi$ and $\chi^{\prime}$ are both $\mathsf{LexiMin}$ optimal.
By the definition of $i$ and the assumption $h_{q}\geq h^{\prime}_{q}+2$, it is easy to see $q>i$ and hence $h_{q}\geq h_{i}$. Further, we claim
$h_{q}\geq h_{i}+2$ (8)
It reduces to prove $h^{\prime}_{q}\geq h_{i}$. Since $\chi$ and $\chi^{\prime}$ are both $\mathsf{LexiMin}$ optimal, the two multisets $\{h_{1},\dots,h_{n}\}$ and $\{h^{\prime}_{1},\dots,h^{\prime}_{n}\}$ are equivalent. By (6), for all $k\in[i-1],h_{k}=h^{\prime}_{k}$. Together, the multiset $\{h_{i},\dots,h_{n}\}$ equals $\{h^{\prime}_{i},\dots,h^{\prime}_{n}\}$. Further by the assumption $h_{1}\leq\dots\leq h_{n}$, the elements in $\{h^{\prime}_{i},\dots,h^{\prime}_{n}\}$ are not smaller than $h_{i}$. Recall $q>i$, we have $h^{\prime}_{q}\geq h_{i}$.
Recall that the family of clean bundles $\mathcal{I}_{j}=\{S\subseteq[m]\mid v_{j}(S)=|S|\}$ for $j\in[n]$ forms a family of independent sets of a matroid Benabbou et al. (2021). By ((I3)) of the independent-set matroid axioms and inequality (7), there exists an item $o_{1}\in\chi^{\prime}_{i}\setminus\chi_{i}$ making $v_{i}(\chi_{i}\cup\{o_{1}\})=v_{i}(\chi_{i})+1=h_{i}+1$. And $o_{1}$ is allocated to some agent $i_{1}\neq i$ under allocation $\chi$. Otherwise, $o_{1}$ is not allocated to anyone, and can be allocated to agent $i$ violating that $\chi$ is max-USW. Consider the following three cases:
  1. (1)
    Suppose $h_{i_{1}}\geq h_{i}+2$. Then transferring $o_{1}$ from $i_{1}$ to $i$ in $\chi$ decreases $\mathsf{LexiMin}(\mathbf{p}(\chi))$, contradicting that $\chi$ is $\mathsf{LexiMin}$ optimal.
  2. (2)
    Suppose $h_{i_{1}}=h_{i}+1$. We note that $i_{1}\neq q$ since $h_{q}\geq h_{i}+2$ (inequality (8)). If we transfer $o_{1}$ from $i_{1}$ to $i$ in $\chi$, $\mathsf{LexiMin}(\mathbf{p}(\chi))$ and $h_{q}$ are unchanged, which means $(\chi,\chi^{\prime})$ still satisfies the two conditions (i) $\chi,\chi^{\prime}\in\mathcal{SS}$ and (ii) $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\geq 2$, but the $\sum_{i\in[n]}|\chi_{i}\triangle\chi^{\prime}_{i}|$ decreases, a contradiction.
  3. (3)
    Suppose $h_{i_{1}}\leq h_{i}$.
    We first show $h_{i_{1}}\leq h^{\prime}_{i_{1}}$. This clearly holds by (6) if $i_{1}<i$.
    When $i_{1}>i$, since $h_{1}\leq\dots\leq h_{n}$, we have $h_{i_{1}}\geq h_{i}$. Together with the assumption $h_{i_{1}}\leq h_{i}$ in this case, we have $h_{i_{1}}=h_{i}$.
    Suppose to the oppose that $h_{i_{1}}>h^{\prime}_{i_{1}}$, which means $h_{i}>h^{\prime}_{i_{1}}$. Applying (6),
    $\qquad\mathsf{LexiMin}((h_{1},\dots,h_{i-1},h_{i}))>\mathsf{LexiMin}((h^{\prime}_{1},\dots,h^{\prime}_{i-1},h^{\prime}_{i_{1}})).$
    Moreover,
    $\mathsf{LexiMin}((h^{\prime}_{1},\dots,h^{\prime}_{i-1},h^{\prime}_{i_{1}}))\geq\mathsf{LexiMin}((h^{\prime}_{j_{1}},\dots,h^{\prime}_{j_{i}})),$
    since $h^{\prime}_{j_{1}},\dots,h^{\prime}_{j_{i}}$ are the smallest $i$ elements in $\mathbf{p}(\chi^{\prime})$. Together, $\mathsf{LexiMin}((h_{1},\dots,h_{i}))>\mathsf{LexiMin}((h^{\prime}_{j_{1}},\dots,h^{\prime}_{j_{i}}))$, by comparing the smallest $i$ elements of $\mathbf{p}(\chi)$ and $\mathbf{p}(\chi)$, yielding $\mathsf{LexiMin}(\mathbf{p}(\chi))>\mathsf{LexiMin}(\mathbf{p}(\chi^{\prime}))$, a contradiction.
    With $h_{i_{1}}\leq h^{\prime}_{i_{1}}$ (i.e. $v_{i_{1}}(\chi_{i_{1}})\leq v_{i_{1}}(\chi^{\prime}_{i_{1}})$) in hand, since $\chi_{i_{1}}$ is clean, we have $v_{i_{1}}(\chi_{i_{1}}\setminus\{o_{1}\})<v_{i_{1}}(\chi^{\prime}_{i_{1}})$. Further since $\chi_{i_{1}}\setminus\{o_{1}\}$ and $\chi^{\prime}_{i_{1}}$ are clean (i.e. independent sets of a matroid), there exists an item $o_{2}\in\chi^{\prime}_{i_{1}}\setminus(\chi_{i_{1}}\setminus\{o_{1}\})$ such that $v_{i_{1}}(\chi_{i_{1}}\setminus\{o_{1}\}\cup\{o_{2}\})=v_{i_{1}}(\chi_{i_{1}})=h_{i_{1}}$. And $o_{2}$ is allocated to some agent $i_{2}\neq i_{1}$ under $\chi$, as otherwise $o_{2}$ is not allocated to anyone and we can transfer $o_{1}$ from $i_{1}$ to $i$ and allocate $o_{2}$ to $i_{1}$ in $\chi$ violating that $\chi$ is max-USW. We note that $\chi^{\prime}_{i_{1}}\setminus(\chi_{i_{1}}\setminus\{o_{1}\})=\chi^{\prime}_{i_{1}}\setminus\chi_{i_{1}}$ and $o_{2}\neq o_{1}$ because $o_{1}\notin\chi^{\prime}_{i_{1}}$ (recall that $o_{1}\in\chi^{\prime}_{i}\setminus\chi_{i}$) and $o_{2}\in\chi^{\prime}_{i_{1}}$.
Repeating the same argument and letting $i_{0}=i$, we obtain a sequence of items and agents $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$. Let $\chi^{(k)}$ denote the allocation that transferring $o_{l}$ from $i_{l}$ to $i_{l-1}$ under $\chi$ for all $l\in[k]$. The sequence of items and agents satisfying $o_{k}\in\chi^{\prime}_{i_{k-1}}\setminus\chi^{(k-1)}_{i_{k-1}}$.
[…]
Figure 3. Transferring items along the sequence $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$. Some items are denoted by $a,\dots,g$.
To sum up, suppose to the opposite of Theorem 22 that there are some pairs of allocations $(\chi,\chi^{\prime})$ that meets (i) $\chi,\chi^{\prime}\in\mathcal{SS}$ and (ii) $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\geq 2$, then it leads to some contradiction. Thus, for $\chi,\chi^{\prime}\in\mathcal{SS}$, it holds that $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\leq 1$. ∎
Remark 0.
Our proof of Theorem 22 is similar to that of Lemma 3.12 in Benabbou et al. (2021), which is the 0/1-sub valuation version of Lemma 6 (implying an allocation non-optimal under $\mathsf{LexiMin}$ is non-optimal under any SPD criterion). There is a minor flaw in their proof of Lemma 3.12 in Benabbou et al. (2021) which we have modified in our proof of Theorem 22.
In the proof of Lemma 3.12 in Benabbou et al. (2021), for the sequence of items and agents $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$, they argue that no same agent appears twice, which implies that if the same agent appears again, the allocation $\chi$ is still clean after transferring items along the cycle (which is a subsequence of the sequence). However, this is not true. See the example in Figure 3, where the sequence is $(3,a,2,b,1,c,2,d,4)$. After transferring items along the cycle $(2,b,1,c,2)$ (which is equivalent to swapping items $b$ and $c$ between agents $2$ and $1$), the allocation may not be clean if $\{a,b\}$ is not a clean bundle of agent $2$. Indeed, the family of clean bundles of agent $2$ may be $\mathcal{I}_{2}=\{\{a,c\},\{a,d\},\{b,c\},\{b,d\},\{a\},\{b\},\{c\},\{d\}\}$ such that $\{a,b\}\notin\mathcal{I}_{2}$.
We note that Lemma 3.12 in Benabbou et al. (2021) still holds and the minor flaw in its proof in Benabbou et al. (2021) can be corrected by arguing that no same item appears twice like ours proof of Theorem 22.
Summary
For the binary valuations, it is known the set of optimums has nothing to do with the exact criterion, as long as the criterion is SPD and symmetric. A complete proof to this result was scattered in literature and is provided in this paper. Furthermore, this paper discusses the consistency among the set of SPD optimal allocations, that is, their profiles are close to each other under Chebyshev distance, which holds for 0/1-sub valuations. The proof is nontrivial.
For the 0/1-add valuations, we introduce a layer partition among items and agents. Together with the idea of dividing items into $2n^{2}$ identical pieces, we obtain the first combinatorial algorithm for computing an optimal allocation (i.e. a stable allocation) for the divisible items, which has a lower complexity than the LP approach.
Appendix A All Criteria in Example 2 are SPD
Claim 1.
Minimizing $\mathsf{EnvySum}(\mathbf{p})$ is equivalent to minimizing $\mathsf{GiniIndex}(\mathbf{p})$.
Proof.
${\sum}_{h_{i}<h_{j}}(h_{j}-h_{i})={\sum}_{i<j}(h^{\uparrow}_{j}-h^{\uparrow}_{i})$
$=$ ${\sum}_{j}(h^{\uparrow}_{j}(j-1))-{\sum}_{i}(h^{\uparrow}_{i}(n-i))$
$=$ ${\sum}_{i}(h^{\uparrow}_{i}(i-1-n+i))$
$=$ $2{\sum}_{i}(i\cdot h^{\uparrow}_{i})-(1+n){\sum}_{i}(h^{\uparrow}_{i})$
$=$ $2\mathsf{GiniIndex}(\mathbf{p})-(1+n)m.$
Lemma 24.
Consider two profiles $\mathbf{p}=(p_{1},\ldots,p_{n})$ and $\mathbf{q}=(q_{1},\ldots,q_{n})$, where $\mathbf{q}$ is more balanced that $\mathbf{p}$ (namely, there are $j,k\in[n]$ such that $p_{j}>q_{j}\geq q_{k}>p_{k}$ and $q_{i}=p_{i}$ for $i\in[n]\setminus\{j,k\}$).
  1. (1)
    $\mathsf{NSW}^{\neg}(\mathbf{q})<\mathsf{NSW}^{\neg}(\mathbf{p})$.
  2. (2)
    $\mathsf{Potential}_{\Phi}(\mathbf{q})<\mathsf{Potential}_{\Phi}(\mathbf{p})$, where $\Phi(x)$ is a strictly convex function of $x$. (Hence criteria $\sum_{i}\Phi(h_{i})$ with strictly convex terms $\Phi(h_{i})$, including $\mathsf{Entropy}^{\neg}(\mathbf{p})$, $\mathsf{Congestion}(\mathbf{p})$, $\mathsf{LexiMax}(\mathbf{p})$, and $\mathsf{LexiMin}(\mathbf{p})$ are all SPD.)
  3. (3)
    $\mathsf{EnvySum}(\mathbf{q})<\mathsf{EnvySum}(\mathbf{p})$.
    (Hence $\mathsf{GiniIndex}(\mathbf{q})<\mathsf{GiniIndex}(\mathbf{p})$ due to 1.)
Therefore all criteria mentioned in Example 2 are SPD.
Proof.
By definition, we assume $q_{j}=p_{j}-\Delta$ and $q_{k}=p_{k}+\Delta$ where $\Delta>0$. Then we have $p_{j}-p_{k}\geq 2\Delta$.
1.
$\mathsf{NSW}^{\neg}(\mathbf{q})-\mathsf{NSW}^{\neg}(\mathbf{p})$
$=$ $-(q_{j}q_{k}-p_{j}p_{k})\prod_{i}{p_{i}}$
$=$ $(p_{j}p_{k}-(p_{j}-\Delta)(p_{k}+\Delta))\prod_{i}{p_{i}}$
$=$ $(p_{j}p_{k}-p_{j}p_{k}-\Delta p_{j}+\Delta p_{k}+\Delta^{2})\prod_{i}{p_{i}}$
$=$ $(\Delta(\Delta-(p_{j}-p_{k}))\prod_{i}{p_{i}}$
$\leq$ $(\Delta(\Delta-2\Delta))\prod_{i}{p_{i}}$
$=$ $-(\Delta^{2})\prod_{i}{p_{i}}<0.$
Thus we have $\mathsf{NSW}^{\neg}(\mathbf{q})<\mathsf{NSW}^{\neg}(\mathbf{p})$.
2.
$\mathsf{Potential}_{\Phi}(\mathbf{q})-\mathsf{Potential}_{\Phi}(\mathbf{p})$
$=$ $(\Phi(q_{k})-\Phi(p_{k}))+(\Phi(q_{j})-\Phi(p_{j}))$
$=$ $(\Phi(p_{k}+\Delta)-\Phi(p_{k}))-(\Phi(q_{j}+\Delta)-\Phi(q_{j})).$
Since $\Phi(x)$ is a strictly convex function of $x$, and $q_{j}>p_{k}$, we have
$(\Phi(p_{k}+\Delta)-\Phi(p_{k}))<(\Phi(q_{j}+\Delta)-\Phi(q_{j})).$
Thus,
$\mathsf{Potential}_{\Phi}(\mathbf{q})<\mathsf{Potential}_{\Phi}(\mathbf{p}).$
3. Let $n_{1},n_{2},n_{3},n_{4}$ and $n_{5}$ be the number of agents $i$ except $j,k$ with $p_{i}\leq p_{k}$, $p_{k}<p_{i}\leq q_{k}$, $q_{k}<p_{i}<q_{j}$, $q_{j}\leq p_{i}<p_{j}$ and $p_{i}\geq p_{j}$ respectively.
Let $U(\mathbf{p},i)={\sum}_{\begin{subarray}{c}j\\ h_{i}<h_{j}\end{subarray}}(h_{j}-h_{i})$ denote the unhappiness of agent $i$. Then we have
$\mathsf{EnvySum}(\mathbf{p})={\sum}_{i}U(\mathbf{p},i),$
$\mathsf{EnvySum}(\mathbf{q})-\mathsf{EnvySum}(\mathbf{p})={\sum}_{i}(U(\mathbf{q},i)-U(\mathbf{p},i)).$
Observe the change from $\mathbf{p}$ to $\mathbf{q}$, we derive that:
  • (i)
    For those agents $i$ with $p_{i}\leq p_{k}$ or $p_{i}\geq p_{j}$, their $U(\mathbf{p},i)=U(\mathbf{q},i)$s remain unchanged.
  • (ii)
    For a agent $i$ with $p_{k}<p_{i}\leq q_{k}$, its $U(\mathbf{p},i)$ changes by
    $U(\mathbf{q},i)-U(\mathbf{p},i)=(q_{k}-p_{i}+q_{j}-p_{i})-(p_{j}-p_{i})=q_{k}-p_{i}-\Delta\leq 0.$
  • (iii)
    For those agents $i$ with $q_{k}<p_{i}<q_{j}$, their total sum of unhappiness is changed by
    $\Delta_{3}={\sum}_{{\begin{subarray}{c}i\\ q_{k}<p_{i}<q_{j}\end{subarray}}}(U(\mathbf{q},i)-U(\mathbf{p},i))=-n_{3}\Delta.$
  • (iv)
    For a agent $i$ with $q_{j}\leq p_{i}<p_{j}$, its unhappiness changes by
    $U(\mathbf{q},i)-U(\mathbf{p},i)=-(p_{j}-p_{i})\leq 0.$
  • (v)
    For $j$, its unhappiness changes by
    $\Delta_{j}=U(\mathbf{q},j)-U(\mathbf{p},j)\leq(n_{4}+n_{5})\Delta,$
    because of the decrease from $p_{j}$ to $q_{j}$.
  • vi
    For $k$, its unhappiness changes by
    $\Delta_{k}=U(\mathbf{q},k)-U(\mathbf{p},k)\leq-(n_{3}+n_{4}+n_{5})\Delta-2\Delta.$
    The $(n_{3}+n_{4}+n_{5})\Delta$ part is due to the increase from $p_{k}$ to $q_{k}$ and the remaining $2\Delta$ part is due to the increase from $p_{k}$ to $q_{k}$ and the decrease from $p_{j}$ to $q_{j}$.
Since case (i), (ii) and (iv) contribute non-positive terms in $\mathsf{EnvySum}(\mathbf{q})-\mathsf{EnvySum}(\mathbf{p})$, we may only consider (iii), (v) and (vi) and have:
$\mathsf{EnvySum}(\mathbf{q})-\mathsf{EnvySum}(\mathbf{p})$
$=$ ${\sum}_{i}(U(\mathbf{q},i)-U(\mathbf{p},i))$
$\leq$ $\Delta_{3}+\Delta_{j}+\Delta_{k}$
$\leq$ $-n_{3}\Delta+(n_{4}+n_{5})\Delta-(n_{3}+n_{4}+n_{5})\Delta-2\Delta$
$=$ $-n_{3}\Delta-2\Delta<0.$
Thus we have $\mathsf{EnvySum}(\mathbf{q})-\mathsf{EnvySum}(\mathbf{p})<0$. ∎
Appendix B Proofs of Lemma 6 and Theorem 9
Proof of Lemma 6.
Assume $\chi$ is non-optimal under $\mathsf{LexiMin}$. We shall prove that $\chi$ is non-stable, i.e., it admits a narrowing transfer.
First, take an allocation $\chi^{*}$ that is optimal under $\mathsf{LexiMin}$.
We build a graph $G$ with $n$ vertices. If an item is allocated to $j$ in $\chi$ and allocated to $k$ in $\chi^{*}$, where $k\neq j$, build an arc from $j$ to $k$. Note that $G$ may have duplicate arcs. Clearly, an arc represents a reallocation (of one item) on $\chi$, and $\chi$ becomes $\chi^{*}$ after all the arcs (i.e. reallocations) are applied.
We decompose $G$ into several cycles $C_{1},\ldots,C_{a}$ and paths $P_{1},\ldots,P_{b}$. Denote by $s_{i},t_{i}$ the starting and ending vertices of $P_{i}$, respectively. We assume that $t_{j}\neq s_{k}$ for any $j\neq k$; otherwise we connect the two paths $P_{j},P_{k}$ into one path.
For $0\leq i\leq b$, let $\chi^{(i)}$ be the allocation copied from $\chi$ but applied all the arcs (reallocations) in $P_{1},\ldots,P_{i}$. $\chi^{(0)}=\chi$.
Note that $\chi^{(b)}$ becomes $\chi^{*}$ after applying the arcs in $C_{1},\ldots,C_{a}$. We obtain that $\mathsf{LexiMin}(\chi^{(b)})=\mathsf{LexiMin}(\chi^{*})$. Further since that $\mathsf{LexiMin}(\chi^{*})<\mathsf{LexiMin}(\chi)$, there exists $i\penalty 10000\ (1\leq i\leq b)$ such that $\mathsf{LexiMin}(\chi^{(i)})<\mathsf{LexiMin}(\chi^{(i-1)})$.
It follows that in $\chi^{(i-1)}$, we have $h_{s}\geq h_{t}+2$ (where $s,t$ denote $s_{i},t_{i}$ respectively, for short). It further follows that in $\chi^{(0)}=\chi$, we also have $h_{s}\geq h_{t}+2$, as $h_{s}$ never increases and $h_{t}$ never decreases in the sequence $\chi^{(0)},\ldots,\chi^{(i-1)}$. Consequently, $\chi$ admits a narrowing transfer (from $s$ to $t$). ∎
Proof of Theorem 9.
Suppose to the opposite that $h_{i}\geq h^{\prime}_{i}+2$. We shall prove that $\chi$ admits a narrowing transfer and hence is nonstable, which contradicts the assumption $\chi\in\mathcal{S}$.
We build a graph $G$ with $n$ vertices. If an item is allocated to $j$ in $\chi$ and allocated to $k$ in $\chi^{\prime}$, where $k\neq j$, build an arc from $j$ to $k$. Note that $G$ may have duplicate arcs. Clearly, $\chi$ becomes $\chi^{\prime}$ after all the arcs (i.e. reallocations) are applied.
Decompose $G$ into paths $P_{1},\ldots,P_{b}$ and a few cycles. Denote by $s_{j},t_{j}$ the starting and ending vertices of $P_{j}$, respectively. We assume that $t_{j}\neq s_{k}$ for $j\neq k$ as in Lemma 6. Moreover, assume that $s_{1}=s_{2}=i$ (due to $h_{i}\geq h^{\prime}_{i}+2$).
For $0\leq j\leq b$, let $\chi^{(j)}$ be the allocation copied from $\chi$ but applied all the arcs (reallocations) in $P_{1},\ldots,P_{j}$. $\chi^{(0)}=\chi$.
Because $\mathbf{p}(\chi^{(b)})=\mathbf{p}(\chi^{\prime})$, we have $\mathsf{LexiMin}(\chi^{b})=\mathsf{LexiMin}(\chi^{\prime})$. We also know $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{\prime})$ since both $\chi^{(0)},\chi^{\prime}$ are stable (Lemma 6). Together, $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(b)})$. It further implies that $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(1)})=\ldots=\mathsf{LexiMin}(\chi^{b})$. Otherwise, there exists $j$ such that $\mathsf{LexiMin}(\chi^{(j)})$ ¡ $\mathsf{LexiMin}(\chi^{(j-1)})$, which means $\chi^{(j-1)}$ admits a narrowing transfer $s_{j}\rightarrow t_{j}$, implying that $\chi^{(0)}$ admits a narrowing transfer $s_{j}\rightarrow t_{j}$.
Since $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(1)})=\mathsf{LexiMin}(\chi^{(2)})$, we know $s_{1}\rightarrow t_{1}$, i.e. $i\rightarrow t_{1}$ (recall $s_{1}=i$), is a swapping transfer of $\chi^{(0)}$, whereas $s_{2}\rightarrow t_{2}$, i.e. $i\rightarrow t_{2}$ (recall $s_{2}=i$), is a swapping transfer of $\chi^{(1)}$. It follows that $h_{t_{2}}=h_{i}-2$, therefore $\chi^{(0)}=\chi$ admits a narrowing transfer $i\rightarrow t_{2}$. ∎
Appendix C Find a stable allocation for indivisible items in $O(m^{2}n)$ time
According to Theorem 7 Claim 1, finding a stable allocation reduces to finding the allocation with minimum $\mathsf{Congestion}$.
Obviously, the latter further reduces to computing the minimum-cost flow in the following network (see Figure 4):
There are $m+n+2$ nodes, including a source node $s$, a sink node $t$, and $m$ nodes $u_{1},\ldots,u_{m}$ representing the items and $n$ nodes $v_{1},\ldots,v_{n}$ representing the agents. And there are $\Theta(mn)$ edges in the network: (i) an edge from $s$ to each $u_{i}$, with capacity 1 and cost 0; (ii) an edge from $u_{i}$ to $v_{j}$ if agent $j$ likes item $i$, with capacity 1 and cost 0; (iii) $m$ edges from each $v_{j}$ to $t$, in which the $k$-th one has capacity 1 and cost k1k-1.
Figure 4. Reduction to minimum-cost flow problem.
Our target – a flow of size $m$ with the minimum cost – can be computed by the Successive Shortest Path algorithm, which increases the size of the current flow by 1 via augmenting along the shortest path in the residual graph, repeating $m$ times. For our particular network, finding such a path reduces to finding a non-used edge $(v_{j},t)$ with lowest cost such that $s$ can reach $v_{j}$ in the residual network, which can be done in $O(mn)$ time by BFS. In total it runs in $O(m^{2}n)$ time.
Appendix D Inconsistency for the Mixed Case
When some items are divisible and others are indivisible, the optimal allocations under different SPD criteria may differ. In other words, the consistency on optimums among all the SPD criteria for divisible and indivisible case, respectively, does not generalize to the mixed case. We show an example in the following. Suppose there are 4 agents and 6 items. The first 4 items are indivisible whereas items $5,6$ are divisible. The matrix $a$ below represents the preference of agents: item $i$ can be allocated to agent $j$ if and only if $a_{i,j}=1$:
$a=\left(\begin{array}[]{llll}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 1&0&0&1\\ 0&1&1&1\\ 0&1&1&1\end{array}\right)$
We use a matrix $b$ to represent the allocation, where $b_{i,j}$ indicates the amount of item $i$ allocated to agent $j$. Note that $b_{i,j}=0$ if $a_{i,j}=0$ and the $i$-th row of matrix $b$ contains only one 1 if item $i$ is indivisible. The sum of each row of the matrix equals 1, and the sum of each column equals the amount of items obtained by each agent.
In this example, all allocations can be divided into two classes depending on whether item $4$ is assigned to agent $1$ or agent $4$. It is not hard to verify:
The optimal allocation for minimizing $\mathsf{LexiMin}$ is shown in $b_{\mathsf{LexiMin}}$ below, admitting profile $(2,\frac{4}{3},\frac{4}{3},\frac{4}{3})$.
The optimal allocation for minimizing $\mathsf{LexiMax}$ is shown in $b_{\mathsf{LexiMax}}$ below, admitting profile $(1,\frac{5}{3},\frac{5}{3},\frac{5}{3})$.
$b_{\mathsf{LexiMin}}=\left(\begin{array}[]{llll}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 1&0&0&0\\ 0&\frac{1}{3}&0&\frac{2}{3}\\ 0&0&\frac{1}{3}&\frac{2}{3}\end{array}\right),b_{\mathsf{LexiMax}}=\left(\begin{array}[]{llll}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 0&\frac{1}{3}&\frac{1}{3}&\frac{1}{3}\\ 0&\frac{1}{3}&\frac{1}{3}&\frac{1}{3}\end{array}\right).$